### v3

# Step-by-Step Mastery: Enhancing Soft Constraint Following Ability of Large Language Models

Abstract
It is crucial for large language models (LLMs) to follow instructions that involve multiple constraints. However, it is an unexplored area to enhance LLMs’ ability to follow soft constraints. To bridge the gap, we initially design a pipeline to construct datasets with high-quality outputs automatically. Additionally, to fully utilize the positive and negative samples generated during the data construction process, we choose Direct Preference Optimization (DPO) as the training method. Furthermore, taking into account the difficulty of soft constraints indicated by the number of constraints, we design a curriculum learning training paradigm based on the constraint quantity. We experimentally evaluate the effectiveness of our methods in improving LLMs’ soft constraint following ability and analyze the factors driving the improvements. The datasets and code are publicly available at https://github.com/Rainier-rq/FollowSoftConstraint.
Step-by-Step Mastery: Enhancing Soft Constraint
Following Ability of Large Language Models
Qingyu Ren1,3, Jie Zeng1, Qianyu He1, Jiaqing Liang2†, Yanghua Xiao1† Weikang Zhou3, Zeye Sun3, Fei Yu3 1Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University 2School of Data Science, Fudan University 3Ant Group {qyren24, jzeng23, qyhe21}@m.fudan.edu.cn, {liangjiaqing, shawyh}@fudan.edu.cn
1 Introduction
In the application of LLMs, the instruction following ability is of paramount importance, especially when the instructions involve multiple constraints Lou et al. (2024); Zeng et al. (2023); Zhou et al. (2023a). The capability of LLMs plays a critical role in aligning with human preferences, ensuring the reliability and helpfulness of the models’ outputs Wang et al. (2023a); Song et al. (2024).
Following instructions with soft constraints is imperative for LLMs Jiang et al. (2023b); Qin et al. (2024). Constraints can be categorized into soft and hard constraints. Hard constraints can be explicitly expressed as specific rules and directly verified through programming methods Zhou et al. (2023a); He et al. (2024a). For example, Python can parse JSON data to verify whether it follows specific format constraints. However, instructions in real-world applications often contain semantic-level limitations, which can be categorized as soft constraints. Soft constraints include restrictions related to content Liang et al. (2024); Zhang et al. (2023), specific backgrounds Shanahan et al. (2023); Liu et al. (2023), and the style of expression Sigurgeirsson and King (2024); Mukherjee et al. (2024). They are difficult to verify automatically through programming methods. As shown in Fig. 1, following soft constraints is challenging for LLMs.
Figure 1: In real-world scenarios, user instructions contain many soft constraints, posing challenges for LLMs. We use bold to represent soft constraints.
Figure 2: The framework of our study. We first design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. Then, we propose a DPO training paradigm based on curriculum learning. CL denotes curriculum learning.
Improving LLMs’ soft constraint following ability is under-explored. Much of the existing work focuses on hard constraints, and it is challenging for LLMs to generalize from hard to soft constraints He et al. (2024a). Additionally, existing research focuses on evaluating LLMs’ soft constraint following ability Chen et al. (2024a); Qin et al. (2024) rather than improving the ability. Overall, existing work on soft constraints has following three limitations: First, directly using advanced models to generate outputs for constructing the dataset results in low quality Xu et al. (2023); Chiang et al. (2023). Soft constraints are especially challenging for these models. On FollowBench Jiang et al. (2023b), GPT-4 demonstrates a hard satisfaction rate of merely 74.4%, making the assurance of high-quality outputs difficult. Second, simply using Supervised Fine-Tuning (SFT) Ouyang et al. (2022) to train the model leads to poor efficiency and performance. Slight changes about the soft constraint in the instruction can result in significant variations in the output He et al. (2024a). It is hard to learn these differences just from positive samples. Third, existing work ignores the difficulty of soft constraints indicated by the number of constraints. Many studies show that organizing training data in the order of the curriculum is better than random shuffling Sun et al. (2024); Lee et al. (2024). Therefore, an efficient training paradigm is required.
In this work, we systematically investigate strategies to enhance the ability of LLMs to follow soft constraints, with the framework shown in Fig. 2, including constructing datasets with high-quality outputs and proposing a DPO training paradigm based on curriculum learning. To construct datasets with high-quality outputs, we progressively add constraints to the instructions and incorporate Judger to reorder the outputs based on constraint following. To fully leverage the positive and negative samples generated during Judger reordering, we apply Direct Preference Optimization (DPO) Rafailov et al. (2024) as the training method. By contrasting positive and negative outputs, it can capture how minor soft constraint adjustments affect output preferences. To account for the difficulty of soft constraints indicated by the number of constraints, we propose a training paradigm that constructs curricula based on the number of constraints in the instruction. In this paradigm, the model starts by learning simpler tasks (fewer constraints) and progresses to more complex ones (more constraints). Our methods can improve LLMs’ soft constraint following ability effectively.
Our contributions are summarized as follows: (1) We design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. (2) We introduce a DPO training paradigm that constructs curricula based on the number of constraints. (3) We conduct extensive experiments to validate the effectiveness of our methods and analyze the reasons for the performance improvements.
2 Related Work
Soft Constraint Following Existing research on soft constraint following focuses on evaluating the ability of LLMs by constructing benchmarks Jiang et al. (2023b); Qin et al. (2024). These benchmarks include a variety of fine-grained constraint types Zhang et al. (2024). These constraints can be categorized into several types: (1) Content soft constraints involve restrictions on the scope or depth of the responses Zhou et al. (2023b); Ashok and Poczos (2024). (2) Situation soft constraints refer to the background limitations Wang et al. (2023b); Shao et al. (2023). (3) Style soft constraints limit the manner of expressions Tao et al. (2024); Pu et al. (2024). Some works directly utilize responses generated by GPT-4 to construct datasets Sun et al. (2024); Peng et al. (2023). However, LLMs struggle to follow soft constraints He et al. (2024b), making responses unreliable. Different from these, we construct datasets with high-quality outputs to improve LLMs’ soft constraint following ability.
Curriculum Learning Curriculum learning is a training strategy that mimics the learning process of humans from simpler to more complex tasks Soviany et al. (2022); Wang et al. (2021). Current research on LLMs’ curriculum learning can be categorized into two primary paradigms: (1) Learning Based on Data Difficulty: This approach organizes the training data sequence according to various evaluation metrics. Metrics such as sequence length Pouransari et al. (2024), perplexity Liu et al. (2024) have been employed to guide this process. LLMs can also construct curricula by organizing the training data sequence in a strategic way Ryu et al. (2024). (2) Learning Based on Task Difficulty: This paradigm focuses on changing the training tasks Chen et al. (2024b) or adjusting the training objectives Zhao et al. (2024b); Lee et al. (2024). However, our work organizes the curriculum based on the number of constraints .
3 Method
In this section, we provide a detailed explanation of how to obtain datasets with high-quality outputs and how to leverage the dataset by establishing a curriculum learning training paradigm. The framework is shown in Fig. 2.
3.1 Dataset Construction
We first construct a multi-constraint instruction following dataset. Adding all constraints at once increases the complexity of the instructions rapidly, making it difficult for the model to understand all constraints He et al. (2024a). To addresss this, we propose a progressive construction method, adding one constraint at a time, which allows the model to understand and learn to follow each constraint effectively. Moreover, existing works in dataset construction rely on advanced models to directly generate the outputs Sun et al. (2024). However, even GPT-4 is struggling to follow the instructions with complex soft constraints Jiang et al. (2023b); Qin et al. (2024). To obtain high-quality outputs, we use Judger to reorder the outputs based on the extent of constraint following. Overall, our pipeline consists of two successive steps: Progressive Construction and Judger Reordering.
3.1.1 Progressive Construction
To enable the model to effectively learn how to follow each constraint, we propose a progressive construction method. Specifically, we add only one constraint at a time, enabling the model progressively learn to follow each constraint during the training process.
We begin by collecting seed instructions from three sources. We first collect instructions from Open Assistant Köpf et al. (2024), which includes instructions generated by users interacting with chatbots. We select rank 0 instructions and those from the first turn of conversations. Next, we gather manually created instructions from the Self-Instruct Wang et al. (2022a). The third source is Super-Natural Wang et al. (2022b), from which we select instructions after filtering out tasks with simple outputs. These three sources together provide a total of 1,500 seed instructions, offering a broad range of coverage across diverse tasks.
Subsequently, we construct different types of soft constraints. Initially, we categorize the soft constraints into three types: content, situation, and style. Next, we randomly select constraints for each seed instruction. For the soft constraint type we select, GPT-4o is employed to generate corresponding descriptions. The prompt used to construct soft constraints is detailed in the Appx. A.1. For the hard constraint type, we select the description from a predefined description list.
To obtain multi-constraint instructions, we adopt a progressive construction approach. As shown in Fig. 2, we add only one constraint to the instruction at a time, allowing the model to focus on learning to follow each constraint. This process helps the model gradually adapt to the increasing complexity of the constraint following task and balance multiple constraints effectively. Specifically, for seed instruction $I_{0}$, we progressively add one constraint each time to form the instruction set $I=\{I_{1},I_{2},\dots,I_{n}\}$, where $n$ denotes the maximum number of constraints. For each instruction $I_{k}$ with $k$ constraints $(k=1,2,\dots,n)$, we use GPT-4o to generate the corresponding output $O_{k}=\text{LLM}(I_{k})$. After performing inference on all the instructions in the instruction set $I$, we obtain the output set $O=\{O_{1},O_{2},\dots,O_{n}\}$.
3.1.2 Judger Reordering
In §3.1.1, we progressively increase the constraints, but the quality of the outputs may not improve incrementally. To address this, we introduce Judger to reorder the outputs based on constraint following to ensure the quality of outputs.
During the progressive construction process in §3.1.1, as new constraints are progressively added, the model’s responses may overlook previously added constraints, leading to a decrease in the output quality. As shown in Fig. 2, to obtain high-quality outputs, we introduce Judger, where GPT-4o is prompted to compare two outputs before and after adding the new constraint, to determine which better follows the updated instruction. The two outputs in each comparison are recorded, and the one deemed better by Judger is used for the next round of comparison. By iteratively ranking the outputs, the constructed data is consistent with constraint following, thereby improving the output quality.
Specifically, when a new constraint is added into the instruction $I_{k-1}$ to form $I_{k}$ , the model’s response $O_{k}$ may not fully follow the constraints in $I_{k}$. To obtain high-quality outputs, we use Judger to rank the new output $O_{k}$ with the previous output $O_{w_{k-1}}$ that more follows $I_{k-1}$ to determine which one better follows the current instruction $I_{k}$: $O_{w_{k}},O_{l_{k}}=\text{Judger}\ (I_{k},O_{w_{k-1}},O_{k}).$
In each ranking, we can obtain the output $O_{w_{k}}$ which follows the current instruction $I_{k}$ better and the output $O_{l_{k}}$ which follows less. Finally, after completing all $n$ rankings, we obtain the positive set $O_{w}=\{O_{w_{1}},O_{w_{2}},\dots,O_{w_{n}}\}$, which consists of outputs that follow their respective instructions better. We also obtain the negative set $O_{l}=\{O_{l_{1}},O_{l_{2}},\dots,O_{l_{n}}\}$, which contains outputs that less follow. The prompt used to reorder outputs and reordering cases are detailed in the Appx. A.2.
3.2 Curriculum Learning Training Paradigm
In §3.1.2, we use Judger to obtain the positive set $O_{w}$ and the negative set $O_{l}$. As shown in Fig. 2, to fully leverage both the positive and negative sets, we apply DPO as the training method. To account for the difficulty of soft constraints indicated by the number of constraints , we establish a curriculum learning training paradigm based on the number of constraints in the instruction.
Given the positive set and the negative set, we can construct the training dataset with $n$ triplets: ($I_{1}$, $O_{w_{1}}$, $O_{l_{1}}$), ($I_{2}$, $O_{w_{2}}$, $O_{l_{2}}$), …, ($I_{n}$, $O_{w_{n}}$, $O_{l_{n}}$). In each triplet, the output from $O_{w}$ is preferred than the output from $O_{l}$. To fully leverage these samples, we apply DPO to train the model. To enable the model to learn from easy to hard tasks during training, we propose a curriculum learning training paradigm based on the number of constraints, starting with simpler tasks involving fewer soft constraints and progressively advancing to more complex tasks involving more soft constraints.
Specifically, for curriculum $k$, the training dataset $D_{k}$ contains the triplet $(I_{k},O_{w_{k}},O_{l_{k}})$, where $I_{k}$ represents the instruction with $k$ constraints, and $O_{w_{k}}$ and $O_{l_{k}}$ denote the corresponding outputs in preference learning. The model begins with the simplest dataset, $D_{1}$, and sequentially progresses through $D_{2}$ to $D_{n}$, gradually enhancing its ability to handle more soft constraints. The complete curriculum is defined as $D=\{D_{1},D_{2},\dots,D_{n}\}$. This stepwise approach ensures that the model first builds a strong foundation by learning from simpler datasets with fewer soft constraints, and gradually adapts to more complex datasets with more constraints. To prevent catastrophic forgetting during training McCloskey and Cohen (1989), we mix each curriculum’s data with a proportion of 10k ShareGPT examples based on its data size Chiang et al. (2023). Based on the curriculum learning training paradigm, the loss function of DPO training is as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-{\mathbb{E}}_{(I_{k},O_{w_{k}},O_{l_{k}})\sim{D}_{k}}[\text{log}\sigma(\beta\text{log}\frac{\pi_{\theta}(O_{w_{k}}|I_{k})}{\pi_{\text{ref}}(O_{w_{k}}|I_{k})}$
$-\beta\text{log}\frac{\pi_{\theta}(O_{l_{k}}|I_{k})}{\pi_{\text{ref}}(O_{l_{k}}|I_{k})})].$
where $\pi_{\theta}$ represents the current model, and $\pi_{\text{ref}}$ denotes the reference model.
To ensure training stability Xu et al. (2024), we add the SFT loss into the DPO loss function:
$\mathcal{L}_{\text{Ours}}=\mathcal{L}_{\text{DPO}}+\mathcal{L}_{\text{SFT}}.$
where SFT loss is as follows:
$\mathcal{L}_{\text{SFT}}(\pi_{\theta})=-\mathbb{E}_{(I_{k},O_{w_{k}})\sim D_{k}}[\log\pi_{\theta}(O_{w_{k}}|I_{k})].$
3.3 Analysis and Comparison
Curriculum # Constraints # Preference Pairs Avg Length
Curri.1 1 3714 369
Curri.2 2 3494 422
Curri.3 3 3387 461
Curri.4 4 3300 503
Curri.5 5 3148 516
Table 1: Statistics of curricula. # Constraints refers to the number of constraints in each instruction. # Preference Pairs refers to the number of preference pairs. “Avg Length” denotes the average instruction length.
Figure 3: The verb frequency in the instructions.
3.3.1 Data Statistics
We present a statistical analysis of different curricula in Tab. 1. The results show that the number of constraints and instruction length in the curriculum continuously increase. Each curriculum contains a large scale of preference data. To show the diversity of our dataset, we analyze the frequency of verbs in the instructions. As shown in Fig. 3, the instructions contain a variety of verbs, reflecting diverse linguistic patterns. This diversity is crucial for enhancing the model’s ability to generalize across different types of constraints.
Method Nums. Cons. Pair. Prog.
Conifer Sun et al. (2024) 13606 H/S $\times$ $\checkmark$
Suri Pham et al. (2024) 20000 S $\checkmark$ $\times$
AutoIF Dong et al. (2024) - H $\checkmark$ $\times$
Complex to Simple He et al. (2024a) 12939 H $\checkmark$ $\times$
Ours 17043 H/S $\checkmark$ $\checkmark$
Table 2: A detailed comparison of related works. Ours represents our dataset. “Nums.”, “Cons.”, “Pair.”, and “Prog.”  denote the number of preference pairs, constraint types, whether to perform pairwise comparison, and whether the dataset is progressively constructed.
3.3.2 Comparison with Other Works
As shown in Tab. 2, we compare our dataset with related works. Our dataset is large in scale compared to these methods. In terms of constraint categories, it includes both soft and hard constraints, enhancing the model’s ability to learn to follow different types of constraints. Additionally, we use Judger for pairwise comparisons of outputs, improving the overall quality of the dataset. Moreover, our dataset is progressively constructed.
4 Experiments
Model BaseModel FollowBench (HSR) IFEval
L1 L2 L3 L4 L5 Avg [S]P [S]I [L]P [L]I Avg
GPT-4 Achiam et al. (2023) GPT 84.7 75.6 70.8 73.9 61.9 73.4 76.9 83.6 79.3 85.4 81.3
GPT-3.5 Turbo GPT 80.3 68.0 68.6 61.1 53.2 66.2 - - - - -
[…]
Qwen2-72B-Instruct Yang et al. (2024) Qwen 82.4 68.2 62.6 61.2 54.7 65.8 77.1 84.4 80.4 86.9 82.2
[…]
Vicuna-13B-v1.5 Chiang et al. (2023) LLaMA2 56.2 42.9 32.3 32.1 24.6 37.6 43.1 53.6 46.6 58.0 50.3
Mistral-7B-ShareGPT Jiang et al. (2023a) Mistral 51.6 45.8 38.3 25.8 20.7 36.5 43.6 53.5 47.3 57.8 50.6
LLaMA3-8B-ShareGPT Dubey et al. (2024) LLaMA3 57.3 49.0 42.3 30.5 27.8 41.4 37.7 49.5 42.3 53.8 45.8
Conifer-7B Sun et al. (2024) Mistral 54.3 49.5 49.3 40.8 30.5 44.9 45.8 57.1 50.8 62.0 53.9
[…]
Mistral-7B-Instruct-v0.3SFT+Judger Mistral 58.7 52.4 42.5 37.2 35.6 45.3 56.8 67.8 60.6 71.3 64.1
Mistral-7B-Instruct-v0.3DPO+Judger+CL Mistral 63.3 56.0 47.5 40.0 36.7 48.7 53.4 63.5 57.1 67.5 60.4
[…]
LLaMA2-13B-Chat-HFSFT+Judger LLaMA2 53.6 43.0 41.2 34.9 25.8 39.7 32.3 44.7 43.1 55.0 43.8
LLaMA2-13B-Chat-HFDPO+Judger+CL LLaMA2 53.6 48.3 42.5 29.2 29.9 40.7 35.1 46.6 48.2 58.9 47.2
[…]
LLaMA3-8B-InstructSFT+Judger LLaMA3 66.3 55.4 50.1 49.7 39.8 52.3 70.4 77.8 73.2 80.1 75.4
LLaMA3-8B-InstructDPO+Judger+CL LLaMA3 69.2 59.6 50.8 48.9 44.6 54.6 72.5 80.3 77.1 84.1 78.5
Table 3: The overall performance on FollowBench and IFEval. We use bold for the best results and underlined for the second-best results among the models ranging from 7B to 13B parameter sizes. indicates that the results are directly sourced from the original benchmarks.
We conduct extensive experiments to validate the effectiveness of our method on improving LLMs’ soft constraint following ability.
4.1 Experiment Setup
Models. We conduct experiments on three widely recognized base LLMs, LLaMA3-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), and LLaMA2-13B-Chat-HF Touvron et al. (2023). Within our experimental framework, we compare three approaches: (1) BASE directly utilizes the base model to generate outputs. (2) SFT+Judger applies supervised fine-tuning on LLMs using the instruction-response pairs $(I_{n},O_{w_{n}})$ generated by Judger (§3.1.13.1.2). (3) DPO+Judger+CL utilizes Judger to obtain high-quality training data, which is then used for DPO training following the curriculum learning paradigm (§3.1.2, §3.2).
For baseline comparisons, we select various open-source and proprietary LLMs. Among the proprietary models, we include GPT-4 Achiam et al. (2023) and GPT-3.5 Turbo. Additionally, we compare our approach with several open-source LLMs, including models specifically trained to improve general instruction following abilities and models focusing on enhancing the ability to follow complex instructions. We also compare our models with two 70B-sized powerful models.
Settings.
For each seed instruction, we progressively add five constraints. In the curriculum setting, we combine curriculum 1 to 3 into a simple curriculum and curriculum 4 to 5 into a difficult curriculum. The model is first trained on the simple curriculum, and then on the difficult one.
Evaluation Benchmarks.
IFEval Zhou et al. (2023a) is a benchmark designed to evaluate the ability to follow hard constraints. It defines 25 distinct types of verifiable instructions, each containing between 1 and 3 constraints. FollowBench Jiang et al. (2023b) is a benchmark that evaluates the ability of models to follow both soft and hard constraints across multiple levels of granularity.
4.2 Main Results
As shown in Tab. 4, our method significantly enhances the model’s ability to follow soft constraints. Compared to the BASE method, there is a significant performance improvement across both benchmarks when three models are trained using the DPO+Judger+CL method, particularly on IFEval. The improvement is significant on difficult soft constraint following tasks, especially at the L5 difficulty level in FollowBench. Specifically, LLaMA-3-8B-Instruct shows an improvement of 5.5% at the L5 difficulty level.
After training with the DPO+Judger+CL method, within the range of models with 8B parameters, Mistral-7B-Instruct-v0.3 and LLaMA-3-8B-Instruct outperform other models on both benchmarks. Within the 13B model category, LLaMA2-13B-Chat-HF outperforms WizardLM-v1.2-13B and Vicuna-13B-v1.5 on FollowBench. Although its performance on IFEval, which only includes hard constraints, is lower than that of the other two models, our method demonstrates superior results on FollowBench, a more complex task that incorporates both soft and hard constraints.
After training with the SFT+Judger method, there is a decline in the performance of LLaMA2-13B-Chat-HF and Mistral-7B-Instruct-v0.3. This drop is attributed to the model’s integration of various specialized training techniques during its initial training phase. Although the SFT+Judger method performs better than DPO+Judger+CL on IFEval for the Mistral-7B-Instruct-v0.3 model, the method even performs worse than the BASE method on FollowBench, failing to effectively improve the model’s ability to follow soft constraints. Moreover, the performance of the SFT+Judger method is worse than the DPO+Judger+CL method on both benchmarks for the other two models.
4.3 Generalization Experiments
Model BaseModel AlpacaEval2.0 MT-Bench
GPT-4-0613 GPT 30.2 9.18
GPT-3.5-Turbo-0613 GPT 22.4 8.39
[…]
WizardLM-13B-v1.2 LLaMA2 14.5 7.20
Vicuna-13B-v1.5 LLaMA2 10.5 6.57
[…]
DPO+Judger+CL LLaMA3 22.0 6.80
Table 4: Results of the length control win rate of AlpacaEval2.0 and the score of MT-Bench. indicates that the results are directly sourced from the original leaderboards.
Besides the ability to follow soft constraints, we also assess the model’s general instruction following abilities on AlpacaEval2.0 Zhao et al. (2024a) and MT-Bench Zheng et al. (2023). We first perform SFT on the model, followed by DPO+Judger+CL. Specifically, we use precomputed outputs of GPT-4 Turbo on AlpacaEval as reference outputs and GPT-4o as evaluators. As shown in the Tab. 4.3, our method improves the model’s general instruction following ability on both banchmarks.
4.4 Ablation Studies
Model FollowBench (HSR) IFEval
L1 - L3 L4 - L5 Avg Avg
BASE 56.3 44.9 51.7 74.3
SFT 59.5 38.4 51.0 73.8
SFT+Judger 57.3 44.8 52.3 75.4
DPO+Judger 58.8 44.6 53.1 78.4
DPO+Judger+CL 59.9 46.8 54.6 78.5
Table 5: Ablation study results on FollowBench and IFEval.
Curriculum Setting FollowBench (HSR) IFEval
L1 - L3 L4 - L5 Avg Avg
BASE 56.3 44.9 51.7 74.3
{C1}{C2}{C3}{C4}{C5} 57.4 37.9 49.6 76.7
{C1,C2,C3}{C4,C5} 59.9 46.8 54.6 78.5
Table 6: Results of different curriculum setting.
In this section, we conduct ablation experiments to study the impact of Judger and curriculum learning on LLaMA3-8B-Instruct. As shown in Tab. 5, directly using the constructed data for SFT without Judger reordering underperforms the SFT+Judger method on both benchmarks, even worse than the base model. The performance decreases significantly at the L4-L5 levels of FollowBench. This suggests that Judger plays a critical role in ranking the model’s responses to challenging instructions. The model trained with DPO outperforms the SFT baseline, especially on IFEval, emphasizing the effectiveness of DPO over SFT in constraint following tasks. Additionally, the curriculum learning training paradigm improves the model’s ability to follow constraints on both benchmarks, particularly those at higher difficulty levels (L4-L5) of FollowBench. This validates the necessity of curriculum learning paradigm for enhancing the model’s ability to follow soft constraints.
We also explore the impact of curriculum setting. As shown in Tab. 6, combining Curriculum 1 to Curriculum 3 into an easy curriculum and Curriculum 4 to Curriculum 5 into a difficult curriculum significantly improves model’s performance on both benchmarks compared to five separate curricula, especially on FollowBench. If the model is trained on five separate curricula, there is no clear boundary between task complexities. The model struggles to generalize from easier tasks to more difficult ones. Grouping similar tasks together with a clear progression in difficulty helps the model better focus on mastering each level of complexity, preventing overfitting to individual tasks.
4.5 Analysis
Figure 4: Results across various constraint categories.
4.5.1 Category Analysis
In this section, we analyze the model’s performance across different types of constraints. Specifically, we compare the performance of LLaMA3-8B-Instruct with the BASE and DPO+Judger+CL method on FollowBench. As shown in Fig. 4, our method significantly improves the model’s performance across different types of constraints. The most notable improvement is observed in the Mixed category , which is defined as a composition of multiple constraint categories to simulate complex real-world scenarios. Our method enhances the model’s performance by 10.6% in this category, suggesting a notable enhancement in the model’s ability to handle complex constraints in real-world scenarios. For categories that contain only soft constraints, our method improves the model’s performance by 6% in the Style category. However, in the category of situation constraints, the performance of the DPO+Judger+CL method is worse than that of BASE. The reason is that our constructed situation constraints focus on soft constraints, while in FollowBench, the situation constraints include both soft and hard constraints. Our method struggles with the hard constraints in the situation category, such as complex situation reasoning, leading to the decrease in performance.
4.5.2 The Role of Progressive Construction
In this section, we analyze the role of progressively constructing the dataset. We compare two methods: w/o progressive and w/ progressive. w/o progressive means replacing both the instruction field and the chosen field with the corresponding ones from samples that include all constraints in the progressively constructed DPO dataset. As shown in Tab. 7, our progressive construction method enhances the model’s performance on FollowBench, especially at the difficlut level L4. Introducing all constraints at once makes the training process unstable. By progressively adding constraints, the model can focus on learning to follow one constraint at a time and learn how to balance constraints effectively.
Method FollowBench (HSR)
L1 L2 L3 L4 L5 Avg
BASE 67.8 54.5 46.6 50.6 39.1 51.7
w/o progressive 69.8 56.7 50.8 42.4 47.6 53.5
w/ progressive 69.2 59.6 50.8 48.9 44.6 54.6
Table 7: Results of different construction methods.
Ranking Method Kendall Tau Coefficient Position Consistency
w/o Judger 0.847 0.743
w/ Judger 0.862 0.794
Table 8: Results on Judger’s effectiveness.
Figure 5: Results of the model across different training stages in curriculum learning.
4.5.3 The Role of Judger
In this section, we investigate the role of Judger in obtaining high-quality outputs. Specifically, we randomly select 100 seed instructions and rank the outputs at each step of progressive construction . We evaluate the rankings in three scenarios: (1) w/o Judger: directly use GPT-4o to generate outputs without reordering, (2) w/ Judger: use Judger to reorder the outputs, and (3) rankings annotated by human experts, which serve as the reference standard for comparison.
To assess the similarity between these rankings, we employ two metrics. The first is the Kendall Tau Coefficient Kendall (1938),which measures the correlation between two rankings by assessing the agreement in the order of paired items. Another metric is position consistency, which quantifies the proportion of elements that occupy the same relative positions. As shown in Tab. 8, the rankings adjusted by the Judger exhibit greater alignment with human-annotated rankings. This suggests that Judger enhances the consistency of outputs with human judgments, thereby improving their quality.
4.5.4 The Role of Curriculum Learning
We analyze the effects of the curriculum learning paradigm. Specifically, we compare the performance of LLaMA3-8B-Instruct with the DPO+Judger+CL method across three training stages. Stage0 represents the base model, while Stage3 and Stage5 represent stages where the model completes the easy curriculum and the hard curriculum, respectively.
As shown in Fig. 5, our training paradigm progressively enhances the model’s constraint following capability across various training stages. Specifically, after easy curriculum learning, the model trained in Stage3 shows superior performance compared to the base model across L1-L3. The model’s performance at L4-L5 in Stage3 is lower than Stage0. The reason is Stage3 has not adequately prepared for the complexity of L4-L5. The gap between these difficulty levels leads to the performance drop. Subsequentially, when the model progresses to Stage5, after hard curriculum learning, the average performance improves significantly at the difficlut levels L4-L5. The results on IFEval further support this conclusion. Stage0 has the lowest average performance across all indicators. After curriculum learning, there is a significant improvement in the model’s performance on IFEval. Although Stage 5 performs slightly worse than Stage 3 on IFEval, Stage 5 shows a significant performance improvement on FollowBench, indicating the enhancement in the model’s ability to follow soft constraints.
5 Conclusion
In this paper, we systematically study how to improve LLMs’ ability to follow soft constraints. Initially, we design a pipeline to automate the construction of datasets with high-quality outputs for soft constraint following. Based on the pipeline, we introduce a method utilizing positive and negative samples generated during the pipeline. Moreover, we propose a new training paradigm that leverages curriculum learning to enhance LLMs’ soft constraint following ability. The experiment results show that our methods enhance models’ ability to follow soft constraints effectively.
6 Limitations
We discuss the limitations of our study as follows. First, we improve the model’s ability to follow soft constraints, thereby improving its overall instruction following capability. However, even when the model’s output meets all the specified constraints, it may still struggle to fully comply with complex instructions due to limitations in reasoning ability or the knowledge it masters. Additionally, while the dataset constructed encompasses a diverse set of tasks, it may still not cover some task types in the long tail. We consider these as key directions for future research.
Appendix A Appendix
A.1 Details of Soft Constraints
We utilize GPT-4o to construct soft constraints. The three categories of soft constraints that we define are as follows:
  • Soft Constraints in Content: Content soft constraints refer to limitations associated with the data itself. These constraints govern the elements of information, the logical relationships between them, and the scope of topics that need to be covered in the response. When multiple content soft constraints are imposed, the model is required to not only generate comprehensive and coherent content but also ensure that the response aligns with the specific logical definitions and boundaries outlined by the instruction. This presents a significant challenge, as it demands both the integration of diverse elements and the maintenance of internal consistency. To address this challenge, we define the following tasks for constructing and applying content soft constraints:
    1. 1.
      Inclusion of Key Elements: The response must incorporate the key points specified in the instruction. This requires the model to effectively extract and integrate relevant information, ensuring that the essential components are included without omitting critical details.
    2. 2.
      Topic Focus: The model must narrow the discussion to a specific subtopic, avoiding broad generalizations or irrelevant tangents. This task emphasizes the importance of maintaining focus and precision within the scope defined by the instruction.
    3. 3.
      Strict Structure: The generated content must adhere to a predefined structure, such as being organized into coherent paragraphs, utilizing subheadings, or following a specific format. This task imposes a higher demand on the model’s ability to generate well-organized and structured outputs, aligning with the required presentation structure.
    We provide the prompt template for constructing the Content Soft Constraint in Tab. 10 and Tab. 11.
  • Soft Constraints in Situation: Situation soft constraints are those related to the context within which the response is situated. These constraints require the response to be adjusted according to the context or assumptions specified in the instruction, ensuring that the content is appropriate to the given background. Such adjustments may involve factors like a particular time or location, the assumption of a specific role, or drawing conclusions based on certain premises. The response must dynamically adapt to situational changes and maintain consistency with the contextual elements. The tasks defined by these constraints can be categorized as follows:
    1. 1.
      Role-Playing: The response must be framed from the perspective of a specific role or persona, ensuring alignment with the contextual expectations associated with that role.
    2. 2.
      Decision Support: The response should provide advice or recommendations that support decision-making within a particular context.
    3. 3.
      Storytelling: The response should construct a narrative that is situated within a defined time, location, or background, maintaining coherence with the provided contextual elements.
    We provide the prompt template for constructing the Situation Soft Constraint in Tab. 12, Tab. 13, and Tab. 14.
  • Soft Constraints in Style: Style soft constraints pertain to the mode of expression, encompassing factors such as the formality or informality of tone, the level of conciseness in language, and the emotional tenor. These constraints require the response to adjust its style in accordance with the given requirements, adapting to different linguistic contexts. The following task types are defined under this category:
    1. 1.
      Tone Requirement: The generated content must adopt a specific tone, such as formal, humorous, or otherwise defined.
    2. 2.
      Language Complexity Control: The complexity of the language used must adhere to specific standards, such as maintaining conciseness and clarity or employing academic expressions.
    3. 3.
      Emotional Expression: The response must convey a particular emotion, such as positivity or sadness, as dictated by the context.
    We provide the prompt template for constructing the Style Soft Constraint in Tab. 15.
A.2 Details of Judger Reordering
We utilize GPT-4o to reorder the outputs. We provide the prompt of Judger ranking in Tab. 16 and examples of how the Judger ranks responses in Tab. 9.
Judger Ranking Case
Instruction Please write a brief description of quantum physics, ensuring that the terms "wave function" and "superposition state" are included. Avoid using overly complex formulas. Keep the response under 100 words.Ensure that this description is easy to understand.
Model Responses Output (a) : Quantum physics studies the behavior of tiny particles. The *wave function* describes a particle’s state, while *superposition state* means a particle can exist in multiple states at once.
Output (b) : Quantum physics studies the behavior of tiny particles. In the quantum world, particles don’t have fixed positions or states; they can be in many states at once, and only take a definite state when measured.
Judger Output [[A]]. From the perspective of constraint adherence, Output (a) is better. It clearly includes both required terms, "wave function" and "superposition state," and explains them in a simple and concise way. The response stays under 100 words and avoids complex formulas, making it easy to understand.Output (b), on the other hand, doesn’t mention "wave function" and its explanation of "many states" is less clear, making it less aligned with the instruction.
Table 9: The case for Judger ranking outputs.
You are an Instruction Rewriting Expert. You need to rewrite #Given Instruction# based on #Rewriting Requirement#, in order to obtain a #Rewritten Instruction#. Basically, #Rewritten Instruction# should adhere to the following guidelines: 1. Your rewriting cannot omit the non-text parts such as the table and code in #Given Instruction#. 2. #Rewritten Instruction# must be reasonable and must be understood and responded to by humans. 3. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#. /* The Given Instruction */ {Given Instruction}
/* Rewriting Requirement */
Please add one proper content constraint to the #Given Instruction#. The content constraints include but are not limited to: 1. Add a Subtask or Another Related Question. 2. Narrow Down the Topic: Instead of a general theme or topic, provide a more specific subset. 3. Set a Higher Standard: Raise the bar for what’s considered acceptable or successful. 4. Limit Resources: Restrict the number or type of resources someone can use. 5. Introduce Specific Criteria: Mandate particular components or features that must be included. 6. Specifying Sequence: Dictate the order in which certain steps or actions should be taken. Please output in JSON format with the fields ’modified_instruction’ for the modified instruction and ’added_constraint’ for the added constraint.
Table 10: The prompt template for constructing the open-ended question answering task in Content Soft Constraint Jiang et al. (2023b).
You are an Instruction Rewriting Expert. You need to rewrite #Given Instruction# based on #Rewriting Requirement#, in order to obtain a #Rewritten Instruction#. Basically, #Rewritten Instruction# should adhere to the following guidelines: 1. Your rewriting cannot omit the non-text parts such as the table and code in #Given Instruction#. 2. #Rewritten Instruction# must be reasonable and must be understood and responded to by humans. 3. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#. /* The Given Instruction */ {Given Instruction}
/* Rewriting Requirement */
Please add one proper content constraint to the #Given Instruction#. The content constraints include but are not limited to: 1. Specify Language Complexity: Determine whether the text should use simple, intermediate, or advanced language. 2. Control Output Length: Set limits on the text’s length, such as maximum word count or number of paragraphs. 3. Restrict Vocabulary: Include or exclude specific words or phrases, or limit the range of vocabulary. 4. Mandate Structure: Require a specific format, such as headings, bullet points, or a particular narrative style. Please output in JSON format with the fields ’modified_instruction’ for the modified instruction and ’added_constraint’ for the added constraint.
Table 11: The prompt template for constructing the language limitations in Content Soft Constraint.
You are an Instruction Rewriting Expert. You need to rewrite #Given Instruction# based on #Rewriting Requirement#, in order to obtain a #Rewritten Instruction#. Basically, #Rewritten Instruction# should adhere to the following guidelines: 1. Your rewriting cannot omit the non-text parts such as the table and code in #Given Instruction#. 2. #Rewritten Instruction# must be reasonable and must be understood and responded to by humans. 3. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#. /* The Given Instruction */ {Given Instruction}
/* Rewriting Requirement */
Please add one proper situation constraint to the #Given Instruction#. The situation constraints include but are not limited to: 1. Define the Context: Specify a particular situation or environment that the suggestions should be relevant to. 2. Introduce a Specific Problem: Focus on addressing a distinct problem or challenge that needs suggestions. 3. Impose Urgency: Include a time constraint or urgency for when the suggestions should be applied. 4. Limit Options: Restrict the scope of potential suggestions to a narrower set of choices. 5. Add Dependencies: Require that suggestions consider certain conditions or prerequisites. 6. Prioritize Outcomes: Highlight specific outcomes or goals that the suggestions should aim to achieve. Please output in JSON format with the fields ’modified_instruction’ for the modified instruction and ’added_constraint’ for the added constraint.
Table 12: The prompt template for constructing the suggestion generation task in Situation Soft Constraint.
You are an Instruction Rewriting Expert. You need to rewrite #Given Instruction# based on #Rewriting Requirement#, in order to obtain a #Rewritten Instruction#. Basically, #Rewritten Instruction# should adhere to the following guidelines: 1. Your rewriting cannot omit the non-text parts such as the table and code in #Given Instruction#. 2. #Rewritten Instruction# must be reasonable and must be understood and responded to by humans. 3. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#. /* The Given Instruction */ {Given Instruction}
/* Rewriting Requirement */
Please add one proper situation constraint to the #Given Instruction#. The situation constraints include but are not limited to: 1. Specify a Role: Clearly define the role or persona to be taken on during the role-play. 2. Define the Setting: Outline the environment or context in which the role-play should occur. 3. Add Conflict or Challenge: Introduce a specific problem, conflict, or challenge that must be addressed within the role-play. 4. Limit the Actions: Restrict the types or number of actions that can be taken during the role-play. 5. Set Specific Goals: Define clear objectives that the role-player must achieve. 6. Introduce Time Constraints: Impose a time limit for the role-play to unfold or for certain actions to be completed. Please output in JSON format with the fields ’modified_instruction’ for the modified instruction and ’added_constraint’ for the added constraint.
Table 13: The prompt template for constructing the role-playing task in Situation Soft Constraint.
You are an Instruction Rewriting Expert. You need to rewrite #Given Instruction# based on #Rewriting Requirement#, in order to obtain a #Rewritten Instruction#. Basically, #Rewritten Instruction# should adhere to the following guidelines: 1. Your rewriting cannot omit the non-text parts such as the table and code in #Given Instruction#. 2. #Rewritten Instruction# must be reasonable and must be understood and responded to by humans. 3. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#. /* The Given Instruction */ {Given Instruction}
/* Rewriting Requirement */
Please add one proper situation constraint to the #Given Instruction#. The situation constraints include but are not limited to: 1. Define Character Archetypes: Specify certain archetypes or roles characters should fulfill, such as a hero, mentor, or antagonist. 2. Include Specific Plot Points: Mandate the inclusion of certain events or plot twists that must occur. 3. Moral Dilemmas: Introduce a scenario where the characters must make a tough decision that involves competing ethical principles or risks. Please output in JSON format with the fields ’modified_instruction’ for the modified instruction and ’added_constraint’ for the added constraint.
Table 14: The prompt template for constructing the story generation task in Situation Soft Constraint.
You are an Instruction Rewriting Expert. You need to rewrite #Given Instruction# based on #Rewriting Requirement#, in order to obtain a #Rewritten Instruction#. Basically, #Rewritten Instruction# should adhere to the following guidelines: 1. Your rewriting cannot omit the non-text parts such as the table and code in #Given Instruction#. 2. #Rewritten Instruction# must be reasonable and must be understood and responded to by humans. 3. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#. /* The Given Instruction */ {Given Instruction}
/* Rewriting Requirement */
Please add one proper style constraint to the #Given Instruction#. The style constraints include but are not limited to: 1. Tone and Emotion: Specify the desired emotional tone for the response. 2. Writing Style: Ask the AI to mimic a specific author’s writing style. 3. Contradiction: Ask the AI to provide a response that contradicts the previous statement or take a stance opposite to its prior response. 4. Ambiguity: Instruct the AI to create responses with intentional ambiguity or double meanings. 5. Humor or Satire: Request that the response be humorous or satirical, requiring the AI to generate jokes or witty remarks.
Table 15: The prompt template for constructing the Style Soft Constraint Jiang et al. (2023b).
You are a helpful assistant who reviews a debate between two other assistants in evaluating the quality of the outputs for a given instruction.The two assistants, Assistant (a) and Assistant (b), are given an instruction. Output (a) and Output (b) are generated by two different AI chatbots respectively. Assistant (a) and Assistant (b) have conflicting evaluations. Your goal is to review their evaluations and give your final decision on which output is better. Here are some rules of the evaluation: (1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc. (2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction. (3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better. Output your final verdict by strictly following this format: "[[A]]" if Output (a) is better, "[[B]]" if Output (b) is better, and "[[C]]" for a tie. /* Given instruction */ {question} /* The Start of Output (a) */ {answer of assistant a} /* The End of Output (a) */ /* The Start of Output (b) */ {answer of assistant b} /* The End of Output (b) */
Table 16: The prompt template for Judger to reorder the responses Zheng et al. (2023)
A.3 Implementation Details
We train LLaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and LLaMA2-13B-Chat-HF using LLaMA-Factory Zheng et al. (2024) on 4 NVIDIA A100 80GB GPUs, applying LoRA Hu et al. (2021) for efficient training. The lora target is set to all, with all models training for 3 epochs. The per device train batch size is set to 1, and gradient accumulation steps is set to 8. The warm-up ratio is set to 0.1. For SFT, LLaMA-3-8B-Instruct has a learning rate of 1.0e-4, Mistral-7B-Instruct-v0.3 uses 5.0e-7, and LLaMA2-13B-Chat-HF uses 1.0e-7. For DPO, the learning rate is 5.0e-6 with a beta value of 0.1. We apply cosine learning rate scheduler.